OpenStudy (anonymous):

The x-intercept of a line is -5 and the y-intercept of the line is -2. What is the equation of the line?

OpenStudy (anonymous):

Well, if you know that the line stops at (-5,0) and (0,-2), you can find the slope by taking the change in y over the change in x. The slope is then -2/5. Then, you take the x of one coordinate and put it into the equation, finding what must be added to get y. -2/5*0=0. To get -2, you would need to subtract 2. Therefore, b is -2. So your line, in y=mx+b form, is y=-2/5x-2. Hope this helps!

OpenStudy (anonymous):

it does thanks so much!! i have two more problems i dont get can you help me?

OpenStudy (anonymous):

Sorry, I was eating lunch.

OpenStudy (anonymous):

I can try the other problems

OpenStudy (anonymous):

ok thanjs! i have no clue on this one Write an equation of the line perpendicular to the given line that contains P. y space equals space 1 fifth x space plus space 1; P(2, -7)

OpenStudy (anonymous):

but i think this one Write the equation of the line parallel to the given line containing point P. y space equals space 8 x space minus space 1; P(-6, 2) is y=8x+50

OpenStudy (anonymous):

youre supposed to use point slope from for both!

OpenStudy (anonymous):

Okay, so for the first line you need to find the slope. A perpendicular slope is the reverse, so instead of 1 over 5 it is negative 5 over 1, or -5.

OpenStudy (anonymous):

Do both lines in the first problem contain P??

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is it -5x+3??

OpenStudy (anonymous):

and is the second one right??

OpenStudy (anonymous):

Yep!! Good job!! :)

OpenStudy (anonymous):

for which one??

OpenStudy (anonymous):

so the first and second are both right??

OpenStudy (anonymous):

Both!!

OpenStudy (anonymous):

thanks!!!!!!!!